OpenStudy (anonymous):

differentiate f(x)=(sqroot (x^2+1))(sqroot(6x+5)) everything divided by (sqroot(x^2-1))

OpenStudy (anonymous):

combine them all into the one square root: sqrt[(x^2 + 1)(6x + 5)/(x^2 - 1)] so if we call whatever's inside the square root g, then by the chain rule, the derivative is g'/2sqrt(g) so to find g', you gotta use the product rule and the quotient rule: g' = [[2x(6x + 5) + 6(x^2 + 1)](x^2 - 1) - 2x(x^2 + 1)(6x + 5)]/(x^2 - 1)^2 = 2(3x^4 - 12x^2 - 10x - 3)/(x^2 - 1)^2 so the derivative overall is: (3x^4 - 12x^2 - 10x - 3)/sqrt(g)*(x^2 - 1)^2 = (3x^4 - 12x^2 - 10x - 3)*sqrt(x^2 - 1)/[sqrt(x^2 + 1)*sqrt(6x + 5)*(x^2 - 1)^2]

OpenStudy (anonymous):

guess it can be simplified to: (3x^4 - 12x^2 - 10x - 3)/[sqrt(x^4 - 1)*sqrt(6x + 5)*(x^2 - 1)]